OpenStudy (anonymous):

Which of the values in the set {1, 2, 3, 4} is a solution to the equation 3x + 3 = 15? 1 2 3 4

OpenStudy (micheal1234):

4 is your answer x=4

OpenStudy (texaschic101):

3x + 3 = 15 --- subtract 3 from both sides 3x + 3 - 3 = 15 - 3 3x = 12 -- divide both sides by 3 x = 12/3 x = 4

OpenStudy (anonymous):

thx guys

OpenStudy (texaschic101):

anytime :)